# Deep Learning-based Intrusion Detection Systems: A Survey

### v1

@@ 5.1.1. Offline Graph Summarization
Offline graph summarization requires historical log data to provide global knowledge, which extracts log data from persistent storage, summarizes the data, and pushes back the summarized data to the persistent storage.
In 2011, Xie et al. (Xie et al., 2011) take inspiration from web graphs to summarize provenance graphs. They argue that provenance graphs have similar organizational structure and characteristics to web graphs, such as locality, similarity, and consecutiveness. On this basis, existing web graph compression algorithms, enhanced with name-identified reference list and node-crossing gaps, are applied in provenance graphs. BEEP (Lee et al., 2013a) is developed based on the fact that a long-running execution can be partitioned into individual units. BEEP reverse engineers application binaries and instructions to perform selective logging for unit boundaries and unit dependencies. LogGC (Lee et al., 2013b) is an audit log system equipped with Garbage Collecting capabilities. During the system’s execution, LogGC can be invoked at any time to obtain the current audit logs and then generate audit logs that have been summarized. Xu et al. (Xu et al., 2016) initially put forward an aggregation algorithm PCR that preserves event dependencies during the data reduction process, ensuring high-quality forensic analysis. Subsequently, they propose a summarization algorithm PCAR and utilized domain knowledge to conduct further data summarization. Hossain et al. (Hossain et al., 2018) propose two dependency-preserving summarization approaches, FD and SD. FD is allowed to keep backward and forward forensic analysis results. SD preserves the results of common forensic analysis, which runs backward to find the entry points of intrusions and then runs forward from these points to unveil their impacts. LogApprox (Michael et al., 2020) targets the most space-intensive events found in logs, namely file I/O activity, which can account for up to 90% of the log content. Once a regular expression for a given process is learned, LogApprox matches and eliminates new events that match the regular expression.
@@ 5.1.2. Online Graph Summarization
Online graph summarization performs real-time summarization for continually coming provenance graphs, rather than dealing with a static graph.
ProTracer (Ma et al., 2016) alternates between system event logging and unit-level taint propagation. It has a lightweight kernel module and user space daemon for concurrent, out-of-order event processing. By combining logging and tainting, and building/evaluating a prototype, it boosts provenance tracing. NodeMerge (Tang et al., 2018) is a template-based graph summarization system for online event storage. It can directly work on the system-dependent provenance streams and compress data provenance via read-only file access patterns. Winnower (Hassan et al., 2018) is an extensible audit-based cluster monitoring system. For tasks replicated across nodes in distributed applications, it can define a model over audit logs to concisely summarize the behaviors of multiple nodes, thus eliminating the necessity of transmitting redundant audit records to the central monitoring node. Ma et al. (Ma et al., 2018) propose an online audit log compression system based on kernel cache to achieve high-performance audit logging. It features a multi-layer cache scheme, which is distributed among various kernel data structures and utilizes the cache to detect and compress redundant events. The approach proposed by Zhu et al. (Zhu et al., 2021) includes two real-time graph summarization strategies. The first strategy maintains global semantics, which identifies and removes redundant events that do not affect global dependencies. The second strategy is based on suspicious semantics. SEAL (Fei et al., 2021) is a novel graph summarization approach for causal analysis. Based on information-theoretic observations of system event data, it achieves lossless compression and supports real-time historical event retrieval. FAuST (Inam et al., 2022) is a logging daemon that performs transparent and modular graph summarization directly on system endpoints. FAuST consists of modular parsers that parse different audit log formats to create a unified in-memory provenance graph representation. AudiTrim (Sun et al., 2024) is an online data summarization approach based on event dependency analysis, which can significantly reduce the amount of log data while ensuring the accuracy of intrusion detection.
@@ 5.2. Attack Detection
Attack detection is located at the central position of DL-IDS. The objective of attack detection is to accurately identify malicious system events in log data while minimizing false alarms of normal system behaviors. Based on the types of log data used in DL-IDS, we categorize the attack detection approaches into audit log-based detectors, application log-based detectors, network traffic-based detectors, and cross-log-based detectors.
Table 4. Overview of attack detection approaches.
@@
Data Type Approach Release Time Base Model Use of Graph Detection Style Detection Granularity
Audit Pagoda (Xie et al., 2018) 2018 - Online Edge
Wang et al. (Wang et al., 2018) 2018 LSTM - - Log Sequence
Unicorn (Han et al., 2020) 2020 - Online Node
Shin et al. (Shin and Kim, 2020) 2020 SVM, LR, KNN - - Sequence
DeepHunter (Wei et al., 2021a) 2021 GNN - Node
Fält et al. (Fält et al., 2021) 2021 Transformer - Offline Log Sequence
ANUBIS (Anjum et al., 2022) 2022 BNN - Event Traces
Paradise (Wu et al., 2022) 2022 - Online Node
ShadeWatcher (Zengy et al., 2022) 2022 GNN Offline Node
threaTrace (Wang et al., 2022a) 2022 GraphSAGE Online Node
ProGrapher (Yang et al., 2023b) 2023 - - Subgraph
MAGIC (Jia et al., 2024) 2024 GAT On/Offline Node, Subgraph
Flash (Rehman et al., 2024) 2024 GNN Online Node
Application DeepLog (Du et al., 2017) 2017 LSTM - Online Log Sequence
LogC (Yin et al., 2020) 2020 LSTM - Online Log Sequence
Yu et al. (Yu et al., 2021a) 2021 Bi-LSTM - - Log Sequence
LogST (Zhang et al., 2022) 2022 Sentence-BERT - - Log Sequence
LogBD (Liu et al., 2023) 2023 BERT - - Log Sequence
MDFULog (Li et al., 2023a) 2023 BERT - - Log Sequence
Traffic ACO-SVM (Meng et al., 2020) 2020 SVM - - Sequence
Min et al. (Min et al., 2021) 2021 MemAE - - Network Traffic
SigML++ (Trivedi et al., 2023) 2023 ANN - - Single Log
HDLNIDS (Qazi et al., 2023) 2023 CRNN - - Network Traffic
@@
Audit and Application LogAnomaly (Meng et al., 2019) 2019 LSTM - Online Log Sequence
PROV-GEM (Kapoor et al., 2021) 2021 GCN - Node
LogBERT (Guo et al., 2021) 2021 BERT - - Log Sequence
NeuralLog (Le and Zhang, 2021) 2021 BERT - Online Log Sequence
GAN-EDC (Duan et al., 2021) 2021 GAN - - Single Log
LogLS (Chen et al., 2022) 2022 Dual LSTM - Offline Log Sequence
LogUAD (Wang et al., 2022b) 2022 Word2Vec - Offline Log Sequence
Adanomaly (Qi et al., 2022) 2022 BiGAN - - Single Log
LogBASA (Liao et al., 2023) 2023 BERT - - Log Sequence
LAnoBERT (Lee et al., 2023) 2023 BERT - Online Log Sequence
LogDAPT (Zheng et al., 2023) 2023 BERT - - Log Sequence
ASGNet (Yang et al., 2023a) 2023 RoBERTa - - Log Sequence
TCN-Log2Vec (Jiang et al., 2023) 2023 BERT - - Log Sequence
@@
The overview of attack detection approaches is presented in Table 4 and its corresponding taxonomy is depicted in Figure 6. We note that recent years have also published many research papers for attack detection (Li et al., 2024a, 2023c; Xu et al., 2022; Hassan et al., 2020; Milajerdi et al., 2019; Yu et al., 2021b; Chen et al., 2024; El Khairi et al., 2024; Wang et al., 2023). Yet these papers do not belong to DL-IDS, which are thus excluded hereafter in our survey.
@@ 5.2.1. Audit Log-based Detectors
Audit logs are collected from hosts and provide atomic information to construct data provenance. In terms of learning techniques, detectors based on audit logs can be classified as traditional learning, graph neural network, and sequence neural network.
@@
Traditional learning-based detectors refer to those that utilize fundamental learning techniques. Shin et al. (Shin and Kim, 2020) propose to employ diverse machine learning algorithms to categorize benign and malicious logs. Pagoda (Xie et al., 2018) takes into account the degree of abnormality of individual provenance paths and the entire provenance graph. It can quickly uncover intrusions if a serious compromise has been identified on paths. The detection rate of Pagoda is further improved by considering the behavior representation in the whole provenance graph. Unicorn (Han et al., 2020) is a real-time intrusion detector that efficiently constructs a streaming histogram to represent the history of system executions. The counting results of the histogram are updated immediately if new edges are generated. By exploring the neighborhood relationships of provenance graphs, Unicorn discovers the causal relationships of system entities in the context. Similarly, ANUBIS (Anjum et al., 2022) also employs provenance graphs to capture causal relationships to achieve high detection performance. The core predictive capability of ANUBIS is a Bayesian Neural Network, which can inform users about how confident it is in its predictions. Paradise (Wu et al., 2022) introduces a novel extraction strategy at the system log level to prune and extract process feature vectors, which are then stored in an efficient in-memory database. Paradise is independent of types of operating systems or provenance collection frameworks, and is capable of independently calculating provenance-based dependencies at the detection stage.
[…]
Figure 6. Taxonomy of attack detection approaches.
@@
Graph Neural Network (GNN) is demonstrated to do well in graph-structured DL-IDS (Wei et al., 2021a; Yang et al., 2023b; Jia et al., 2024; Rehman et al., 2024; Zengy et al., 2022). DeepHunter (Wei et al., 2021a) is a GNN-based graph matching approach that is designed with two novel networks: attribute embedding networks that could incorporate Indicators of Compromise (IoC) information and graph embedding networks that could capture the relationships between IOCs. threaTrace (Wang et al., 2022a) emerges as an online approach dedicated to detecting and tracing host-level threats. Its model is custom-tailored on the basis of GraphSAGE (Hamilton et al., 2017), an inductive graph neural network framework, aiming to learn the role of every benign entity within the data source graph. MAGIC (Jia et al., 2024) leverages Graph Attention Network (GAT) (Velickovic et al., 2017) as a crucial constituent of its graph representation module. It employs masked graph representation learning to mimic benign system events and executes multi-granularity detection under diverse supervision levels. MAGIC can be applied in both online detection and offline detection scenarios. MAGIC is able to adapt to concept drift and has minimal computational overhead, which makes it applicable to real-world online APT detection. ProGrapher (Yang et al., 2023b) is a graph embedding-based anomaly detection system that extracts features and detects anomalies by utilizing information in relational networks. ShadeWatcher (Zengy et al., 2022) is a recommendation-guided detector. It models high-order connectivity via item-side information, borrowing the recommendation concepts of user-item interactions and applying them to the security concepts of system entity interactions.
@@
All logs can be formatted as sequential data. Hence, sequence neural networks such as Recurrent Neural Network (RNN) and Transformer are employed in some work (Wang et al., 2018; Fält et al., 2021). The approach of Wang et al. (Wang et al., 2018) employs Long Short-Term Memory (LSTM), a variant of RNN, for anomaly detection. In their work, LSTM is compared with other machine learning algorithms such as Gradient Boosting Decision Tree and Naive Bayes. The results indicate that LSTM can attain the optimal anomaly detection performance when utilizing two feature extraction methods. Fält et al. (Fält et al., 2021) present a log anomaly detection method that is entirely trained based on auxiliary system data and is unsupervised with respect to the target system. It utilizes transformers as its base model, whose performance results reveal great potential of the self-attention mechanism in the field of DL-IDS.
5.2.2. Application Log-based Detectors
Application logs are generated from the installed binaries. Generally, application logs are in the form of natural language text, as the generated templates of them are pre-defined by their developers. Hence, many Natural Language Processing (NLP) techniques are introduced in this type of DL-IDS.
Recurrent Neural Network
As one of the most classic methods, RNN is used commonly in DL-IDS. DeepLog (Du et al., 2017) employs LSTM to model system logs as natural language sequences. As a result, it is able to automatically learn log patterns from normal operations and detect anomalies when there is a deviation between log patterns and the trained model. LogC (Yin et al., 2020) comprises two stages: transforming log messages into log template sequences and component sequences. These two sequences are then input into a combined LSTM model for the detection of abnormal logs. LogC only requires normal log sequences to train the combined model. Both DeepLog and LogC are capable of conducting online detection. Yu et al. (Yu et al., 2021a) propose a novel framework for the automatic detection of log anomalies, which leverages a bidirectional LSTM (Bi-LSTM) model based on an attention mechanism. In contrast to existing log anomaly detection methods, this approach not only pays attention to the sequential and quantitative information of log sequences but also considers the hidden semantic information in logs, thus enhancing the efficiency of anomaly detection.
Transformer Architecture
Transformer (Vaswani et al., 2017) is an effective model for sequence data, which exploits a multi-head self-attention mechanism to realize a large model that can be parallelly computed. The most common Transformer-based model in DL-IDS is BERT (Devlin et al., 2019), which takes advantage of encoding sequence data and pre-training model parameters. LogBD (Liu et al., 2023) utilizes BERT to acquire the semantic information of logs and constructs distance-based anomaly detection via a domain adaptation approach. It employs a Temporal Convolutional Network (TCN) to extract the common characteristics of different system logs and map them into the same hypersphere space. MDFULog (Li et al., 2023a) introduces a BERT-based semantic feature extraction model to preserve the semantics of log messages and map them to log vectors. This effectively eliminates the work randomness and noise injection caused by log template updates and enhances the robustness of anomaly detection. LogST (Zhang et al., 2022) obtains the semantic representation of log events through the SBERT model, taking into account the semantics of each word and the word order relationship, thus facilitating a better understanding of the context of log sequences.

### v2

@@ 5.1.1. Offline Graph Summarization
Offline graph summarization requires historical log data to provide global knowledge, which extracts log data from persistent storage, summarizes the data, and pushes back the summarized data to the persistent storage. In 2011, Xie et al. (Xie et al., 2011) take inspiration from web graphs to summarize provenance graphs. They argue that provenance graphs have similar organizational structure and characteristics to web graphs, such as locality, similarity, and consecutiveness. BEEP (Lee et al., 2013a) is developed based on the fact that a long-running execution can be partitioned into individual units. BEEP reverse engineers application binaries and instructions to perform selective logging for unit boundaries and unit dependencies. LogGC (Lee et al., 2013b) is a summarized audit log system that can be invoked at any time during the system execution. Xu et al. (Xu et al., 2016) propose an aggregation algorithm PCR that preserves event dependencies during log data reduction. They further propose an algorithm named PCAR that utilizes domain knowledge to conduct graph summarization. Hossain et al. (Hossain et al., 2018) propose two dependency-preserving graph summarization approaches, FD and SD. FD is allowed to keep backward and forward forensic analysis results. SD preserves the results of common forensic analysis, which runs backward to find the entry points of intrusions and then runs forward from these points to unveil their impacts. LogApprox (Michael et al., 2020) aims to summarize the most space-intensive events found in logs, namely file I/O activity, which can account for up to 90% of the log content.
@@ 5.1.2. Online Graph Summarization
Online graph summarization performs real-time summarization for continually coming provenance graphs, rather than dealing with a static provenance graph. ProTracer (Ma et al., 2016) alternates between system event logging and unit-level taint propagation. It has a lightweight kernel module and user space daemon for concurrent, out-of-order event processing. NodeMerge (Tang et al., 2018) is a template-based graph summarization system for online event storage. It can directly work on the system-dependent provenance streams and compress data provenance via read-only file access patterns. Winnower (Hassan et al., 2018) is an extensible audit-based cluster monitoring system. For tasks replicated across nodes in distributed applications, it can define a model over audit logs to concisely summarize the behaviors of multiple nodes, thus eliminating the necessity of transmitting redundant audit records to the central monitoring node. The approach proposed by Zhu et al. (Zhu et al., 2021) includes two real-time graph summarization strategies. The first strategy maintains global semantics, which identifies and removes redundant events that do not affect global dependencies. The second strategy is based on suspicious semantics. SEAL (Fei et al., 2021) is a novel graph summarization approach for causal analysis. Based on information-theoretic observations of system event data, it achieves lossless compression and supports real-time historical event retrieval. FAuST (Inam et al., 2022) is a logging daemon that performs transparent and modular graph summarization directly on system endpoints. FAuST consists of modular parsers that parse different audit log formats to create a unified in-memory provenance graph representation. AudiTrim (Sun et al., 2024) is an efficient graph summarization approach that reduces log sizes without impacting user experiences, which allows adaptable deployment on different operating systems.
@@ 5.2. Attack Detection
Attack detection is located at the central position of DL-IDS. The objective of attack detection is to accurately identify malicious system events in log data while minimizing false alarms of normal system behaviors. Based on the types of log data, we categorize the attack detection approaches into audit log-based, application log-based, network traffic-based, and hybrid log-based detectors.
Table 4. Overview and taxonomy of attack detection approaches.
@@
Data Type Taxonomy Approach Release Time Base Model Detection Style Detection Granularity
Traditional StreamSpot (Manzoor et al., 2016) 2018 K-Medoids Online Subgraph
Learning Unicorn (Han et al., 2020) 2020 K-Medoids Online Node, Subgraph
DistDet (Dong et al., 2023) 2023 HST Online Subgraph
Audit ShadeWatcher (Zengy et al., 2022) 2022 TransR Offline Node
Log Graph threaTrace (Wang et al., 2022) 2022 GraphSAGE Online Node
Neural ProGrapher (Yang et al., 2023) 2023 graph2vec Online Subgraph
Network MAGIC (Jia et al., 2024) 2024 GAT Online Node, Subgraph
Flash (Rehman et al., 2024) 2024 GraphSAGE Online Node
R-caid (Goyal et al., 2024) 2024 GNN Offline Node
Argus (Xu et al., 2024b) 2024 MPNN, GRU - Node
Traditional Wei et al. (Xu et al., 2009) 2009 PCA, TF-IDF - Log Entry
Learning Bodik et al. (Bodik et al., 2010) 2010 Logistic Regression Online Log Entry
DeepLog (Du et al., 2017) 2017 LSTM Online Log Entry
LogRobust (Zhang et al., 2019) 2019 Attention LSTM - Log Entry
@@
Application Sequence LogAnomaly (Meng et al., 2019) 2019 template2vec, LSTM Online Log Entry
Log Neural LogC (Yin et al., 2020) 2020 LSTM Online Log Entry
Network NeuralLog (Le and Zhang, 2021) 2021 BERT - Log Entry
PLELog (Yang et al., 2021) 2021 Attention GRU Online Log Entry
SpikeLog (Qi et al., 2023) 2023 DSNN - Log Entry
LogCraft (Zhang et al., 2024) 2024 Meta Learning - Log Entry
Traffic Whisper (Fu et al., 2021) 2021 K-Means - Host
Traditional SigML++ (Trivedi et al., 2023) 2023 ANN - Encrypted Log
Learning OADSD (Zhang et al., 2023a) 2023 Isolation Forest Online Packet
LtRFT (Tang et al., 2023) 2023 LambdaMART Offline Packet
Kitsune (Mirsky et al., 2018) 2018 AutoEncoder Online Packet
MT-FlowFormer (Zhao et al., 2022a) 2022 Transformer - Flow
I2RNN (Song et al., 2022) 2022 I2RNN - Packet
Graph and ERNN (Zhao et al., 2022b) 2022 ERNN - Flow
Sequence Euler (King and Huang, 2023) 2023 GNN, RNN - Flow
Neural pVoxel (Fu et al., 2023) 2023 - - Packet, Flow
Network NetVigil (Hsieh et al., 2024) 2024 E-GraphSage - Flow
Exosphere (Fu et al., 2024) 2024 CNN - Packet
DFNet (Zhao et al., 2024a) 2024 DFNet - Packet
RFH-HELAD (Zhong et al., 2024) 2024 RPGAN, Deep kNN - Packet
Hybrid Hybrid OWAD (Han et al., 2023) 2024 Autoencoder Online Hybrid
FG-CIBGC (Ni et al., 2025) 2025 DisenGCN, ICL - Hybrid
@@
The overview and taxonomy of attack detection approaches are presented in Table 4. We note that recent years have also published many other academic papers for attack detection (Xie et al., 2018; Wu et al., 2022; Li et al., 2023b; Milajerdi et al., 2019a; Yu et al., 2021; Chen et al., 2024; El Khairi et al., 2024; Wang et al., 2023; Hassan et al., 2020a). Yet these papers are slightly related to DL-IDS, which are thus excluded in our survey for conciseness.
@@ 5.2.1. Audit Log-based Detectors
Audit logs are collected from hosts and thus detectors based on them are basically referred to as HIDS. Audit logs provide fine-grained information through provenance graphs to depict system behaviors. Depending on the learning techniques, audit log-based detectors can be further classified as traditional learning and graph neural network.
@@
Traditional learning-based detectors refer to those that utilize naive machine learning techniques. StreamSpot (Manzoor et al., 2016) is a clustering-based anomaly detection that tackles challenges in heterogeneity and streaming nature. Unicorn (Han et al., 2020) is a real-time intrusion detector that efficiently constructs a streaming histogram to represent the history of system executions. The counting results within the histogram are updated immediately if new edges (or events) occur. DistDet (Dong et al., 2023) is a distributed detection system that builds host models in the client side, filters false alarms based on their semantics, and derives global models to complement the host models.
@@
GNN is demonstrated to do well in processing provenance graphs (Yang et al., 2023; Jia et al., 2024; Rehman et al., 2024; Zengy et al., 2022; Wang et al., 2022). ProGrapher (Yang et al., 2023) extracts temporal-ordered provenance graph snapshots from the ingested logs, and applies whole graph embedding and sequence-based learning to capture rich structural properties of them. The key GNN technique leveraged by ProGrapher is graph2vec. ShadeWatcher (Zengy et al., 2022) is a recommendation-guided intrusion detector using provenance graphs. It borrows the recommendation concepts of user-item interactions into security concepts of system entity interactions and analyzes cyber threats in an automated and adaptive manner. threaTrace (Wang et al., 2022) emerges as an online approach dedicated to detecting host-based threats at the node level. Its GNN model is a tailored GraphSAGE (Hamilton et al., 2017) for learning rich contextual information in provenance graphs. MAGIC (Jia et al., 2024) leverages Graph Attention Network (GAT) (Velickovic et al., 2017) as its graph representation module. MAGIC employs masked graph representation learning to incorporate the capability of pre-training. It can adapt to concept drift with minimal computational overhead, making it applicable to real-world online APT detection. Flash (Rehman et al., 2024) is a comprehensive and scalable approach on data provenance graphs to overcome the limitations in accuracy, practicality, and scalability. Flash incorporates a novel adaptation of a GNN-based contextual encoder to encode both local and global graph structures into node embeddings efficiently. R-caid (Goyal et al., 2024) first incorporates root cause analysis into PIDS. Before training GNNs, R-caid links nodes to their root causes to build a new graph, intending to prevent it from mimicry and evasion attacks. Argus (Xu et al., 2024b) finds the performance of the prior IDS is questionable on large scale. It thus devises a form of discrete temporal graph and uses encoder-decoder unsupervised learning to detect different types of attacks.
5.2.2. Application Log-based Detectors
Application logs are generated from the installed binaries. Generally, application logs are in the form of natural language text, namely sequence data. It is thus common to introduce sequence-based DL techniques into application log-based DL-IDS.
Traditional Learning
For traditional learning, Wei et al. (Xu et al., 2009) propose a general methodology to mine rich semantic information in console logs to detect large-scale system problems. Bodik et al. (Bodik et al., 2010) leverage a logistic regression model on a new and efficient representation of a datacenter’s state called fingerprint to detect previously seen performance crises in that datacenter.
@@ 5.2.2. Application Log-based Detectors
Due to the similarity between application logs and natural language texts, sequence neural networks such as Recurrent Neural Network (Hochreiter and Schmidhuber, 1997) and Transformer (Vaswani et al., 2017; Devlin et al., 2019) are widely employed. DeepLog (Du et al., 2017) employs LSTM to model system logs as natural language sequences. It is able to automatically learn benign log patterns and detect anomalies when there is a deviation between log patterns and the trained model. LogRobust (Zhang et al., 2019) finds previous methods do not work well under the close-world assumption and utilizes an attention-based LSTM model to handle unstable log events and sequences. LogAnomaly (Meng et al., 2019) identifies previous studies tend to cause false alarms by using indexes rather than semantics of log templates. Empowered by a novel, simple yet effective method termed template2vec, LogAnomaly is proven to successfully detect both sequential and quantitive log anomalies simultaneously. LogC (Yin et al., 2020) is a new log-based anomaly detection approach with component-aware analysis. It feeds both log template sequences and component sequences to train a combined LSTM model for detecting anomalous logs. NeuralLog (Le and Zhang, 2021) targets the performance caused by log parsing errors such as out-of-vocabulary words and semantic misunderstandings and employ BERT to perform neural representation. PLELog (Yang et al., 2021) is a semi-supervised anomaly detection approach that can get rid of time-consuming manual labeling and incorporate the knowledge on historical anomalies. SpikeLog (Qi et al., 2023) adopts a weakly supervised approach to train an anomaly score model, with the objective of handling a more reasonable premise scenario where a large number of logs are unlabeled. LogCraft (Zhang et al., 2024) is an end-to-end unsupervised log anomaly detection framework based on automated machine learning, which mitigates the cost of understanding datasets and makes multiple attempts for building algorithms.